# Vision Transformers for End-to-End Vision-Based Quadrotor Obstacle Avoidance

### v1

@@ III-C Simulated setup and models
Figure 3: ViT model used in this study. ViT+LSTM uses a multi-layer LSTM placed before the final FC layers.
ViT. We utilize a transformer based encoder inspired by Segformer [segformer], and apply two transformer blocks in hierarchical fashion, followed by upsampling using pixel shuffle [pxshuffle], and mixing of information across hierarchies using a convolution operation. This incorporates information at multiple scales. This ViT is followed by fully connected layers. This very lightweight architecture is presented in Figure 3.
ViT+LSTM. This is identical to the ViT model but includes a multi-layer LSTM before the fully connected layers. ConvNet. A lightweight CNN model that resembles the first successful deep-learning-based approaches for object detection or end-to-end vision-based control for robots. This model is considered a control model, and is the only one significantly smaller than 3M parameters. ResNet-18 [he2016resnet], a large CNN model that learns residuals from the input signal, was also trained and tested but found to perform similarly to the ConvNet model and worse than the UNet model, demonstrating that simply increasing the number of parameters in the ConvNet does not lead to better performance.
UNet+LSTM. The symmetric encoder-decoder architecture popular in segmentation tasks includes skip connections that incorporate local and global information and help prevent vanishing gradients. The reduction to a lower-dimensional feature space also may improve generalizability to unseen environments. A multi-layer LSTM follows the UNet.
Model Trainable parameters CPU (ms) GPU (ms)
ConvNet 235,269 0.195 0.423
LSTMnet 2,949,937 4.75 0.735
UNet+LSTM 2,955,822 5.56 1.67
ViT 3,101,199 5.58 3.83
ViT+LSTM 3,563,663 9.16 1.61
@@
Model UNet ViT Conv LSTM MLP
ConvNet (2) (4)
LSTMnet (2) (2) (3)
UNet+LSTM (2) (2) (3)
ViT (1) (3)
ViT+LSTM (1) (3) (2)
TABLE II: Structure and layer components per model. ConvNet and LSTMnet are structured by the rudimentary layer types themselves. Numbers next to checkmarks indicate how many layers of that type are present in the model.
III-D Hardware setup
We use the Falcon 250 custom quadrotor platform [tao2023seer] (Figure 10(a)) running the kr_mav_control [githubGitHubKumarRoboticskr_mav_control] open-source quadrotor control stack. All experiments are performed in a Vicon motion capture arena for accurate state estimate. The Intel Realsense D435i depth camera utilizes both structured light via an infrared emitter and stereo matching to calculate depth images. These images are resized to match the simulation-size images, then fed directly into the network models. The output velocity command is sent to the control stack, which outputs a SO(3) command (orientation and thrust) that is sent to the PX4 controller [meier2015px4]. While we train models on floating objects in simulation, we test with a single, self-standing object in the real-world.
IV RESULTS
IV-A Collision metrics and success rates
Figure 4: Collision and success metrics for each model, taken over 10 randomized, unseen simulation sphere configuration trials at each forward velocity. (a) Mean number of collisions per trial. Most models do not outperform the expert up to speeds of 5m/s, beyond which the ViT+LSTM model outperforms both the expert and other models. (b) Mean time spent in collision per trial, with each trial lasting between 8.5-20s. This metric tends to increase with collision “severity”, i.e. head-on collisions and large-obstacle collisions. The models outperform the expert, and ViT+LSTM does the best. (c) Success rates, i.e. the fraction of trials containing no collisions. ConvNet and ViT+LSTM are the most consistent models with forward velocity.
We assess models with two metrics quantifying collisions over multiple trials in different, unseen sphere environments in Figure 4; generalization performance is detailed in Section IV-D. Average collision rate describes the number of collisions per trial, and mean time spent in collision counts the time the drone spent in collision with obstacles. Impacts with large obstacles and head-on collisions tend to increase the values of the second metric. Figure 4(a) shows that the average number of collisions per trial generally increases as a function of velocity for every model, as expected due to lower allowed response time and as shown in prior work. However, the ViT+LSTM begins to outperform the expert as speeds increase beyond 5m/s. ConvNet presents similar behavior until 6m/s. As the training dataset included expert trajectories with collisions, it is notable that the ViT+LSTM model outperforms this expert. Furthermore, the models primarily composed of the individual components of the ViT+LSTM model, namely ViT and LSTMnet, do not perform well, while the combination model maintains a significantly lower collision rate as speeds reach 7m/s.
Expert mean time in collision (Figure 4(b)) is particularly bad since each collision caused the waypoint search to fail; however, notably, the poor expert statistics are not learned by any of the models. The ViT+LSTM model outperforms all other models in this metric as well, especially as speeds increase. By observation, the ViT+LSTM model tends to clip the edges of obstacles and have fewer head-on collisions than other models; this results in fewer timesteps spent in collision.
IV-B Path and command characteristics
Figure 5: Visualizing the trajectories taken by each model in a single test environment over multiple trials. (a) Top-down view with $0.1\sigma$ distribution of the mean position, (b) side-view with $0.1\sigma$ distribution of the mean position, (c) 3D perspective of the mean paths.
The trajectory paths in Figure 5 exhibit the paths of each model as the drone flies from the $x=0$ starting point to $x=60$. The ConvNet demonstrates significantly less variation in paths, while the ViT model exhibits changing spread in chosen paths as it progresses through the obstacle field. In contrast, the ViT+LSTM model has a steadily increasing spread as $x$ position increases, possibly suggesting both a balance between model capacity as well as consistency of commands over multiple trials, though this needs further investigation.
Figure 6: Commanded accelerations (a) and computed energy costs (b) for all models in the spheres environment. Lower is considered better for both metrics. The ViT-only model commands high-acceleration values without achieving better performance, but the inclusion of recurrent layers improves this behavior.
Figure 6 presents command characteristics of each model across forward velocities in the spheres environment test set. We observe that smoother velocity commands are more feasible to track by the quadrotor and result in lower commanded acceleration, and the energy cost metric is computed as in [yu2023avoidbench]. The ViT model has high commanded accelerations across all forward velocities, but this is drastically improved with the addition of the LSTM layers for ViT+LSTM; notably this comes without a drop in performance (Figure 4). Generally, models including recurrent (LSTM) components command lower acceleration and have lower energy cost; this is particularly the case in the trees environment (Figure 8(c)) described in Section IV-D.
IV-C Network feature analysis
(a) ConvNet mask.
(b) UNet mask.
(c) ViT mask.
Figure 7: Mask visualizations for each spatial operation used in this comparison study. Color images are shown for viewability, though aligned depth images are used as input to the models. ConvNet generally highlights entire obstacles, while UNet and ViT highlights edges. ViT seems to additionally capture more context around obstacles.
Figure 7 highlights areas of the depth image that return the strongest signal by the convolutional layers in the ConvNet model and UNet model, and that of the attention layers in the ViT model. Convnet highlights full obstacles, with little specificity to their shape. UNet layers specifically highlight the edges and disregard all other areas of the image. The ViT model seems to both capture obstacle edges and surrounding context, likely a result of the known behavior of ViTs to learn relationships between parts of an image i nprior object detection work [vits].
IV-D Generalizability
Figure 8: Generalization to trees: collision and success metrics for each model, taken over 10 randomized tree environments. Since we did not train on trees, we do not compare with the expert. (a) Mean number of collisions per trial. (b) Mean time spent in collision per trial. (c) Success rates, i.e. the fraction of trials containing no collisions.
To test model generalization capabilities, we zero-shot deploy the models in a simulation environment containing realistic looking tree models placed randomly in the scene (“tree environment”). We present collision rates and success rates in Figure 8 that resemble the in-distribution sphere environment testing results, showing that the ViT and ViT+LSTM models generalized well and outperformed other models. As speeds increase to 5m/s, time spent in collision increases in this particular environment, suggesting that collisions are more head-on when they occur for such models.
Figure 9: Demonstration of paths taken by non-ViT and ViT models for the simulated flying through a window task. Only the ViT-based models succeed at this challenging task. One representative example is shown for each category since the models within both categories perform similarly.
We additionally present a zero-shot fly-through-window experiment performed across all models at 3 m/s (Figure 9). While the spheres and trees environment contains obstacles within free space, with multiple viable collision-free paths, this environment is completely in collision except for one viable collision-free path. Only the ViT models (ViT and ViT+LSTM) succeed at this task, and all others fail. This presents additional evidence that ViT models are superior in generalization compared to other state-of-the-art computer vision models, including recurrent models, for end-to-end quadrotor control.
IV-E Ablation against state information
Env Models Velocity (m/s)
3 5 7
Spheres ViT 54 81 27 18 18 18
ViT+LSTM 72 54 45 45 36 9
Convnet 63 27 54 18 36 10
LSTMnet 45 27 9 18 9 9
Unet 27 36 9 18 9 27
Trees ViT 100 90 81 81 80 81
ViT+LSTM 100 90 80 81 70 50
Convnet 100 90 63 81 42 60
LSTMnet 90 90 63 40 63 36
Unet 72 100 63 60 54 45
[…]
We perform an ablation against the included state information of quadrotor orientation and forward velocity by training and testing the models without this data (therefore the only input is the depth image itself), with success rates presented in Table III. This ablation is particularly relevant since flying in the wild may result in noisy state estimates and variable forward velocities. The results show that ViT+LSTM suffered from lack of this information as speeds increased beyond 5m/s, and the ConvNet model suffered across all speeds in the spheres environment. In the tree environment, at low speeds most models either maintained or improved success rates when lacking this data. As speeds increase, ConvNet performs better in this unseen, generalization environment without the data.
IV-F Hardware demonstrations
(a) Falcon 250 custom quadrotor platform with a 402mm tip-to-tip diameter, an Intel Realsense D435 depth camera, a Pixhawk 4 Mini flight controller, and an Intel NUC 10 onboard computer with a i7-10710U CPU. Inference is performed onboard.
(b) ConvNet trial.
(c) ViT+LSTM trial.
Figure 10: Real world experiments evaluating the ConvNet and ViT+LSTM models in a dodging task against a rigid obstacle.
We zero-shot deploy the two most successful and most different models from simulation, ConvNet and ViT+LSTM, onboard the Falcon 250 hardware platform described in Section III-D and observe the path of the drone while it dodges a single object. Figure 10 shows the stark difference in paths taken, which we observed to be consistent over multiple trials. The ConvNet does not begin the dodging maneuver until very late ($<$1m distance to obstacle), and then dodges very closely to the obstacle. The ViT+LSTM, however, begins the dodging maneuver much earlier (2m distance to obstacle) and proceeds to take a wide path around the obstacle. While further analysis is needed, these results suggest that the ViT+LSTM model has benefits over a simple ConvNet-style model for real-world deployment.
V CONCLUSION
We present the use of a vision transformer for end-to-end obstacle avoidance of a quadrotor from depth perception, and compare it to other popular learning-based architectures. All models were trained in the same fashion from a privileged expert via behavior cloning. Models with significant recurrent layers command lower accelerations with lower energy cost but low success rates, while ViT attention-based models without an LSTM layer command high values but have higher success rates. The attention-recurrent (ViT+LSTM) model, however, has desirably low acceleration and energy cost while presenting the best collision and success rates. The attention-based models drastically outperformed all other models in simulated generalization tests of a trees environment and a challenging flying-through-window task. Real world demonstrations show that the attention-recurrent (ViT+LSTM) combination model actuates to dodge much earlier than the convolution-only (ConvNet) model, and analyzing model layer masks suggests that attention layers capture context around obstacles in addition to highlighting the edges of visible obstacles. Further work might explore how sensor noise, poor state estimate, and varying robotic platforms may affect generalization capabilities of these end-to-end models. Additionally, commanding lower-level control commands such as collective-thrust and body-rates (CTBR) might outperform commands of linear velocities when attempting learned, fast and agile flight.
ACKNOWLEDGMENT
We thank members of the Kumar Robotics Lab for their help with hardware experiments, rich research discussions, and generous support. This work was funded by NSF SLES-2331880 and NSF CAREER-2045834.
[…]

### v2

@@ IV-F Ablation against state information
We perform an ablation against providing quadrotor orientation and forward velocity to the models by re-training without this data; success rates are presented in Table II for Spheres and Trees across three velocities. This ablation is particularly relevant to high-speed flight since unstructured or windy environments may cause noisy state estimates and variable velocity. We find that ViT+LSTM performed better with the state information across slow and fast speeds and across environments, while ConvNet only benefited in the Spheres environment but performed substantially worse when tested in the unseen, generalization Trees environment, particularly as speeds increase.
IV-G Real world experiments
(a) Falcon250.
(b) 1m/s trials.
(c) ConvNet vs. ViT+LSTM.
Figure 7: Representative real-world experiments using the Falcon250 (7(a) and Section III-D) evaluating the ConvNet and ViT+LSTM models in the same single, rigid obstacle avoidance task (7(b)). Paths are plotted in 7(c), where lighter shades on the red (ConvNet) and blue (ViT+LSTM) trajectories represent higher lateral velocities. ViT+LSTM consistently avoids earlier than the ConvNet model.
We zero-shot deploy ViT+LSTM in single and multi-obstacle configurations requiring collision avoidance and flight-through-gap behaviors at high speeds up to 7m/s. The model runs inference onboard at 30Hz (CPU) during all trials. Figure 1 shows representative examples of two such trials, with 1(c) showing the depth image input to the model with a red arrow representing the model’s corresponding output velocity command. The quadrotor avoids obstacles in the real world with no model re-training or extensive system fine-tuning; while depth camera images exhibit artifacts not seen in the simulated training data, such as high-frequency patterns, varying background appearance, and propellers in-view (uncropped depth images viewable on project website) our end-to-end learned models performed effectively in many scenes at speeds of 1-7m/s. Figure 7 additionally demonstrates real-world differences in collision avoidance between the ConvNet and ViT+LSTM models.
V CONCLUSION
We present the first use of a vision transformer for end-to-end quadrotor control from depth images, with a comprehensive discussion of collision rates and command characteristics in simulation, comparison to other learned architectures and traditional modular methods, as well as deployment in challenging real-world trials. We find that the attention-based models (ViT and ViT+LSTM) outperform other models at high speeds in in-distribution simulation environments, and greatly outperform other architectures across all speeds in generalization environments, even extending from obstacle avoidance in forward flight to a fly-through-window task where other architectures fail. The best model combines attention and recurrence (ViT+LSTM), and consistently presents low collision rates while also maintaining the lowest energy cost of all models at speeds ranging 3-7m/s. Ablations show that across various speeds and environments, the attention-recurrent model improves with the inclusion of quadrotor state information. We also found that, in contrast to the presented end-to-end architecture, modular baseline methods fail to maintain low collision rates when scaling to varying flight speeds. We provide additional experiments and visualizations, as well as all code, datasets, and pretrained weights on the project website (Section I) to enable further research of attention-based end-to-end robotics control.
ACKNOWLEDGMENTS
This work was funded by NSF SLES-2331880 and NSF CAREER-2045834.